# Multiple Choice Learning of Low Rank Adapters for Language Modeling

### v1

@@ E.4 Additional results
Table 9 presents the results of LoRA-MCL and LoRA-MLE. Almost all the metrics show the same trend in which LoRA-MCL outperforms LoRA-MLE with Beam Search (BS) and Diverse Beam Search (DBS) in terms of quality, although DBS produces more varied outputs. Depending on the rank ($r=8$ or $r=24$), we found that setting $\lambda=1.0$ or $\lambda=0.8$, respectively, yields the best quality scores for DBS. Similar to our experiments on Audio Captioning, increasing the rank of LoRA-MLE results in a slight degradation in performance but improves diversity. Additionally, increasing the number of beams in BS with LoRA-MLE results in slightly improved performance but reduced diversity. In contrast, increasing the number of beams in DBS with LoRA-MLE (with fixed $\lambda$) leads to declines in both quality and diversity. Interestingly, with LoRA-MCL, increasing the number of beams enhances both performance and diversity here.
E.4.1 Artificial multilingual dataset creation
@@
To evaluate the behaviour of LoRA-MCL under a multi-modal distribution, we simulated an artificial bi-modal dataset by automatically translating half of the captions from English to French using T5-small [57], while keeping the prompts in English. More specifically, we randomly sampled half of the images and translated their five associated captions. All the training parameters are the same as those in the experiments on the original TextCaps dataset, except the learning rate, which we set at $2\times 10^{-5}$ (as the maximum value in the scheduler) in both the LoRA-MLE and LoRA-MCL.
@@
E.5 Computation details
@@
We run the experiments mostly on H100 NVIDIA GPUs with $80$ GB of RAM. Trainings and inferences were launched on a single GPU for the Audio Captioning experiments, and up to $8$ GPUs for the Image Captioning experiments. The total computing resources used for this project, including failed experiments, amount to approximately $20{,}000$ GPU hours.

### v2

@@ E.7.2 Additional results
The full results presented in Table 12 show that LoRA-MCL tends to outperform LoRA-MLE with Beam Search (BS) and Diverse Beam Search (DBS) in terms of quality, although DBS produces more varied outputs. Depending on the rank ($r=8$ or $r=24$), we found that setting $\lambda$ to $1.0$ or $\lambda=0.8$, respectively, yields the best quality scores for DBS. Similar to our experiments on Audio Captioning, increasing the rank of LoRA-MLE results in a slight degradation in performance while improving diversity. Additionally, increasing the number of beams in BS with LoRA-MLE results in slightly improved performance but reduced diversity. In contrast, increasing the number of beams in DBS with LoRA-MLE (with $\lambda$ fixed) leads to declines in both quality and diversity. Interestingly, with LoRA-MCL, increasing the number of beams enhances both performance and diversity here.
E.7.3 Artificial multilingual dataset creation
@@
To evaluate the behaviour of LoRA-MCL under a multi-modal distribution, we simulated an artificial bi-modal dataset by automatically translating half of the captions from English to French using T5-small (Raffel et al., 2020), while keeping the prompts in English. More specifically, we randomly sampled half of the images and translated their five associated captions. All the training parameters are the same as those in the experiments on the original TextCaps dataset, except the learning rate, which we set at $2\times 10^{-5}$ (as the maximum value in the scheduler) in both the LoRA-MLE and LoRA-MCL.
@@
E.7.4 Specialization of the hypotheses
Understanding what each hypotheses learn is difficult. The founding paper of MCL by Lee et al. (2016) already introduced the notion of specialization of the hypotheses in the context of classification (see Lee et al. (2016) in Figure 4). We provide additional insights with the French/English controlled experiment of Section 6, where we observed some “unsupervised” specialization of the hypotheses; one of the hypothesis learned English, and the other learned French. We tried to visualize the hypotheses predictions to understand what the hypotheses have learned in Figure 6.
To quantify specialization, we embed every generated caption on the evaluation set (for hypotheses $k\in\{1,\dots,K\}$ and each example $i\in\{1,\dots,N\}$) using a Sentence-BERT model (StyleDistance (Patel et al., 2025)). We trained a linear Support Vector Machine (SVM) (Cortes and Vapnik, 1995) on this space to predict which head produced each caption, using a $70/30$ train-test split on captions. In the French/English experiment, the SVM achieves $99.8\%$ test accuracy for the two-head LoRA-MCL model of Section 6, confirming specialization. Embeddings of the hypotheses are visualized using a Principal Components Analysis (PCA) with two components in Figure 6. Outside this controlled setup, we repeated this analysis on captions from the LoRA-MCL trained on AudioCaps with $K=5$. The SVM reaches $63.7\%$ test-accuracy, over three times better than random choice, indicating a clear specialization of the heads. In contrast, applying the same procedure to captions from the LoRA-MLE baseline (with DBS decoding, $\lambda=0.8$, and beam size = 5) yielded $23.2\%$ accuracy on the test set after SVM fitting, close to random choice and showing no evidence of specialization, likewise for LoRA-MoE (yielding $19.8\%$ accuracy in the same setup). We believe that further understanding of how specialization emerges represents a promising direction for future work.
E.7.5 Qualitative Examples
In this section, we show some qualitative examples (image–predicted caption pairs) that highlight the behavior of LoRA-MCL compared to a baseline model (LoRA-MLE with $r=24$). When the scene contains numerous objects or pieces of information, LoRA-MCL (through its different heads) covers a wider variety of descriptions than diverse beam search decoding. We performed inference with LoRA-MCL using greedy decoding, while we used diverse beam search ($\lambda=0.8$ and using $3$ beams) for the baseline model.
LoRA-MLE. LoRA-MCL.
{A display of Hugo Victor products on a table.} {A display of Hugo Victor products is set up in a store.}
{A display for Hug Victor with a sign that says Hug Victor.} {A store display for Hugo Victor with a shelf full of items.}
{a store display for Hugo Victor with many items} {A store display for Hugo Victor features a variety of plates, bowls, and books.}
LoRA-MLE. LoRA-MCL.
{a person getting food from a vending machine with the number 680 on it} {A girl is putting food into a vending machine that says FEEDO.}
{A woman is getting food from a vending machine that says "Fresh" on it.} {A woman is putting money into a vending machine that says 100.}
{A girl puts a sandwich into a vending machine that says "Fresh" on it.} {A girl is getting a hot dog from a vending machine.}
LoRA-MLE. LoRA-MCL.
{ A Scotch brand glue stick is on a table.} {A bottle of Scotch tape sits on a table.}
{A can of Scotch branded glue is on a table.} {A bottle of Scotch tape sits on a table next to a piece of paper.}
{a scotch tape that is on the ground} {A Scotch tape is on the table next to a piece of paper.}
E.8 Diverse Machine Translation
E.8.1 Experimental setup
We fine-tuned ALMA-7B (based on LLaMA-2-7B (Touvron et al., 2023)) for two epochs on the human-written data collected by Xu et al. (2024a), which includes the WMT’17–WMT’20 test sets and the development and test sets from Flores-200 (Costa-Jussà et al., 2022). Our fine-tuning setup follows Xu et al. (2024a), except that we employ bf16 precision instead of fp16. We use $r=\tfrac{\alpha}{2}=16$, a maximum sequence length of $512$, an effective batch size of $256$, a warm-up ratio of $0.01$ and a peak learning rate of $2\times 10^{-5}$ with an inverse square-root scheduler.
We follow the quality–diversity evaluation protocol of Shen et al. (2019), computing BLEU scores with the SacreBLEU (Post, 2018) library, which uses the commonly used Moses 13a Tokenizer (standard in Machine Translation) instead of the Penn Treebank Tokenizer we used for Audio and Image Captioning. The reported metrics are Leave-One-Out BLEU (Loo-BLEU) and Pairwise-BLEU (see Section 4 of Shen et al. (2019)), using the official codebase. Leave-One-Out BLEU (denoted simply as BLEU in (Shen et al., 2019)) measures the overall quality of the hypothesis set by computing the corpus-level BLEU for each hypothesis against the reference set (higher is better). Pairwise-BLEU considers only the predicted hypotheses, computing BLEU scores for each pair of candidates (lower indicates greater diversity). We define hereafter formally these two metrics, as done in (Shen et al., 2019).
Let $\operatorname{BLEU}\left\{\left(\left[x^{1},\cdots,x^{R}\right],\hat{x}^{k}\right)\right\}_{\{k\in[\![1,K]\!]\}}$ denote the corpus-level BLEU when hypothesis $\hat{x}^{k}$ is evaluated against references $x_{1},\cdots,x_{R}$ as references for each example in the test-set and for each $k\in[\![1,K]\!]$. We also define $x^{-r}]=\{x^{1},\dots,x^{r-1},x^{r+1},\dots,x^{R}\}$.
Leave-One-Out BLEU. Corresponds to an average of each corpus-level BLEU when using hypothesis $k$ against all possible sets of M1M-1 references
$\mathrm{Loo{-}BLEU}=\frac{1}{R}\sum_{r=1}^{R}\operatorname{BLEU}\left\{\left(\left[x^{-r}\right],\hat{x}^{k}\right)\right\}_{\{k\in[\![1,K]\!]\}}\;.$
Pairwise-BLEU. Similar to $\mathrm{mBLEU}_{4}$, Pairwise-BLEU considers only the set of candidates as
$\mathrm{Pairwise{-}BLEU}=\operatorname{BLEU}\left\{\left(\left[\hat{x}^{j}\right],\hat{x}^{k}\right)\right\}_{\{j\in[\![1,K]\!],k\in[\![1,K]\!],j\neq k\}}\;.$
We used it instead of $\mathrm{mBLEU}_{4}$ for consistency with the machine translation community.
E.8.2 Qualitative Examples
We provide some qualitative examples of the predictions on newstest2014. Here, LoRA-MCL uses $\varepsilon=0.05$, $B=3$ and $K=3$, and LoRA-MLE ($r=48$, DBS with $B=6$ and $\lambda=0.8$).
[…]
  • Wie Reuters im Juli erstmals berichtete, ist das Sitzlayout die treibende Kraft hinter der Auseinandersetzung um die neuen Jets.
  • Wie Reuters im Juli erstmals berichtete, ist die Sitzanordnung genau das, was den Kampf zwischen den neuesten Jets antreibt.
  • Wie Reuters erstmals im Juli berichtete, ist die Sitzanordnung genau das, was den Kampf zwischen den neuesten Jets antreibt.
  • Wie Reuters im Juli berichtete, ist es genau das Thema der Sitzanordnung, das den Kampf zwischen den neuesten Flugzeugmodellen antreibt.
  • Wie Reuters erstmals im Juli berichtet hatte, ist das Sitzkonzept genau der Punkt, der das Ringen zwischen den neuesten Jets beleb.
  • Wie Reuters bereits erstmals im Juli berichtete, ist es genau das Layout der Sitze, was den Konkurrenzkampf zwischen den neuesten Jets anschürt.
  • Wie Reuters erstmals im Juli berichtete, ist es das Sitz @-@ Layout, das den Wettbewerb zwischen den aktuellen Flugzeugen antreib.
  • Wie Reuters zuerst im Januar berichtete, ist der Aufbau der Sitzreihen genau das, was den Kampf zwischen den neuesten Jets antreibt.
LoRA-MLE.
Wie Reuters zuerst im Juli berichtete, ist die Sitzanordnung genau das, was den Kampf zwischen den neuesten Jets antreibt .}\}
[…]
LoRA-MCL.
Als Reuters im Juli berichtete, ist die Sitzanordnung genau das, was den Kampf zwischen den neuesten Jets ausmacht.}\}
[…]
Example 2. Input english sentence: $\{$“I was vocal about looking into a whole bunch of things, and this was one of them,”Daley told the paper.}\} References:
  • “Ich äußerte eine ganze Reihe von Dingen, die man in Erwägung ziehen sollte, und das war eines davon”, erklärte Daley dem Magazin. Ich habe mich freimütig darüber geäußert eine ganze Reihe an Sachen ausprobieren zu wollen und das war eine davon, teilte Daley der Zeitung mit.
  • “Ich war lautstark dabei, in eine ganze Reihe von Dingen zu schauen, und das war eines von ihnen”, Daley sagte der Zeitung.
  • “Ich äußerte mich lautstark darüber, dass ich mir eine ganze Reihe von Dingen ansehen wollte. Und das war eines davon", sagte Daley der Zeitung..
  • “Ich verkündete lautstark, mir verschiedene Dinge ansehen zu wollen, und dies war eines davon”Daley hat der Zeitung davon erzählt.
  • “Ich war entschlossen, mich mit einer ganzen Reihe von Dingen zu befassen, und dies war eins davon”, sagte Daley der Zeitung.
  • “Ich habe mündlich geäußert, mich mit einer ganzen Menge Dinge befasst zu haben und das hier gehörte dazu”, erzählte Daley der Zeitung.
  • “Ich habe es deutlich gemacht, dass ich eine ganze Reihe von Sachen unter die Lupe nehmen werde, und das war eine von ihnen”, erzählte Daley der Zeitung.
  • “Ich habe mich dazu geäußert, mir mehrere Dinge ansehen zu wollen, und das war eines davon”, sagte Daley der Zeitung.
  • “Ich hatte geäußert, eine ganze Reihe an Dingen zu untersuchen, und dies war eine von ihnen”, sagte Daley der Zeitung.
LoRA-MLE.
„Ich war laut darüber, dass wir uns mit einer ganzen Reihe von Dingen beschäftigen sollten, und dies war eine davon“, sagte Daley der Zeitung.}\}
$\{$„Ich war laut darüber, dass wir uns mit einer ganzen Reihe von Dingen beschäftigen sollten, und das war eine davon“, sagte Daley der Zeitung.}\}
$\{$“Ich habe mich laut darüber geäußert, dass ich mich in eine ganze Reihe von Dingen einlesen wollte, und das war eine davon”, sagte Daley der Zeitung.}\}
LoRA-MCL.
“Ich war laut darüber, dass ich mich in eine ganze Reihe von Dingen einlesen wollte, und das war eine davon,” Daley sagte der Zeitung.}\}
$\{$“Ich war laut über die Untersuchung einer ganzen Reihe von Dingen, und dies war eines von ihnen”, sagte Daley der Zeitung.}\}
[…]
E.9 Computation details
@@
We run the experiments mostly on H100 NVIDIA GPUs with $80$ GB of RAM. Training and inference were launched on a single GPU for the Audio Captioning experiments, and up to $8$ GPUs for the Image Captioning experiments. The total computing resources used for this project, including failed experiments, amount to approximately 25,00025{,}000 GPU hours.
E.10 Use of Large Language Models
We used LLMs assistants in the preparation of this work. They helped to polish the writing (improving clarity, grammar, and style without altering the content) and serving as a coding assistant (visualization, debugging).